1. A hollow cylinder has an outer diameter of 4cm and an inside diameter of 3.85 cm and outer length of 4.85 cm 

a. What is the thickness in cm? 

b. What is the outer volume in cm3? 

c. What is the inner volume in cm3? 

d. What is the volume of the metal cylinder in cm3? 

 


1
Expert's answer
2021-02-25T11:23:28-0500

Given:

"D=4.00\:\rm cm"

"d=3.85\:\rm cm"

"L=4.85\:\rm cm"

a. The thickness


"h=D-d=4.00-3.85=0.15\:\rm cm"

b. The outer volume


"V_{ou}=\frac{4}{3}\pi D^2h\\
=\frac{4}{3}*3.14*4.00^2*4.85=325\:\rm cm^3"

c. The inner volume


"V_{in}=\frac{4}{3}\pi d^2h\\
=\frac{4}{3}*3.14*3.85^2*4.85=301\:\rm cm^3"

d. The volume of the metal cylinder


"V=V_{ou}-V_{in}=325-301=24\:\rm cm^3"